# CTISum: A New Benchmark Dataset For Cyber Threat Intelligence Summarization

### v1

@@ 3.2. Data Statistics
3.3. Fleiss Kappa Analysis
To verify the consistency of the final summary generation quality, we select 5 experts to execute a consistency evaluation of 50 sampled summaries using Fleiss Kappa, where the scores are rated from 1 to 5, with 1 very negative, 2 negative, 3 neutral, 4 positive, and 5 very positive. As shown in the Table 2, the final Fleiss Kappa score is 0.37, indicating relatively low agreement. This can be attributed to the fact that summary generation belongs to language generation task, where different experts tend to have subjective views on summary quality (e.g., examples that annotated as positive or very positive are all good, but they are inconsistent), which differs greatly from the objectivity of classification tasks. However, we also analyze the score distribution of all experts, finding that summaries rated 3 or above account for 96.4% (241/250) of the total, demonstrating that the quality of the annotated dataset is controllable and accurate. Moreover, we have conduct another round of Fleiss Kappa Analysis on the previous annotation scores from multiple raters. Specifically, we combine score 1 and 2 into one bin, keep score 3 as a separate bin, and combine score 4 and 5 into another bin. This result in a Fleiss Kappa score of 0.613184, which indicate the consistency evaluation.
Table 3. Performance of automatic evaluation on CTIS and APS tasks. LLAMA2 is in the zero-shot setting because of the lack of resources. The best results are highlighted in bold.
@@ 3.3. Fleiss Kappa Analysis
BertSumExt (Liu and Lapata, 2019) 82.15 26.91 5.91 13.09 81.36 20.43 3.52 11.76
MatchSum (Zhong et al., 2020) 83.91 39.31 13.39 20.60 83.75 33.76 9.34 18.91
@@
Transformer (Vaswani et al., 2017) 45.83 34.15 12.05 19.88 45.17 32.94 10.37 18.55
T5 (Raffel et al., 2020) 70.48 43.75 16.58 28.06 69.21 42.41 14.35 27.32
BART-base (Lewis et al., 2020) 70.36 44.45 16.26 27.24 69.24 42.48 14.30 26.23
BART-large (Lewis et al., 2020) 71.21 47.11 18.32 29.20 70.41 45.76 16.88 29.03
@@
Longformer(Beltagy et al., 2020) 71.06 46.97 17.21 28.22 70.31 45.39 15.66 27.09
LLAMA2 (Touvron et al., 2023) (zero-shot) 33.87 21.35 5.93 20.51 32.15 20.89 5.66 19.22
ChatGPT (zero-shot) 65.82 40.18 13.22 24.33 64.31 38.70 12.71 22.98
@@
BertSumExt (Liu and Lapata, 2019) 81.38 20.67 3.06 11.87 81.38 20.35 3.46 11.73
MatchSum (Zhong et al., 2020) 83.93 33.86 9.86 19.17 83.75 33.73 9.35 18.88
@@
Transformer (Vaswani et al., 2017) 40.43 28.22 7.86 16.50 39.15 27.03 6.21 15.48
T5 (Raffel et al., 2020) 68.41 36.20 10.27 24.70 66.06 32.09 7.62 21.78
BART-base (Lewis et al., 2020) 70.47 40.47 12.33 25.28 70.70 39.42 11.13 24.45
BART-large (Lewis et al., 2020) 71.31 41.88 12.76 26.54 70.32 40.25 11.65 25.06
@@
Longformer(Beltagy et al., 2020) 70.98 41.35 11.77 25.36 70.41 39.91 9.78 23.93
LLAMA2 (Touvron et al., 2023) (zero-shot) 21.06 15.68 2.23 13.14 20.38 14.66 2.02 12.28
ChatGPT (zero-shot) 48.77 32.06 9.11 18.69 45.39 30.55 8.61 17.75
4. Problem Formulation
5. Experiments
5.1. Evaluation Metrics
Following papers (Hsu et al., 2022; Mukherjee et al., 2022), BERTScore (Zhang et al., 2020) and ROUGE-$n$ (R-$n$) (Lin, 2004) are taken as evaluation metrics, which are widely used for evaluating the quality of summarization. BERTScore is a metric for evaluating text generation models, which measures the similarity between the generated text and reference text using contextual embeddings from pre-trained BERT. ROUGE-$n$ refers to the overlap of n-grams between the generated and reference summaries. Specifically, ROUGE-$1$ evaluates unigram overlap, ROUGE-$2$ measures bigram overlap, and ROUGE-$L$ calculates longest common subsequence overlap. These metrics compare matching units such as words and phrases between the generated and reference summaries, with higher scores indicating better summarization quality. Among the ROUGE metrics, ROUGE-$L$ generally corresponds best with human judgments of summary quality.
In previous studies, human evaluation is usually conducted by crowdsourcing workers who rate responses on a scale from 1 to 5 from the aspects of correctness, relevancy, etc. However, the criteria can vary widely between different individuals. Therefore, this study adopts the human A/B evaluation for a high inter-annotator agreement. Given the generated summaries of two models A and B, three analysts are prompted to go through an entire CTI report and choose the better one for each of the 80 randomly sub-sampled test instances. For objectivity, annotators include those with and without background knowledge (task-related). The final results are determined by majority voting. If the three annotators reach different conclusions, the fourth annotator will be brought in. We adopt the same human evaluation with paper (Mukherjee et al., 2022): 1) Factual Correctness: which summary can be supported by the source CTI report? 2) Relevance: which summary captures pertinent information relative to the CTI report? 3) Coverage: which summary contains the greatest coverage of relevant content in the CTI report?
5.2. Experimental Setting
@@
The implementation of baselines is based on the HuggingFace framework. The AdamW optimizer (Loshchilov and Hutter, 2017) with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ is used for training, with an initial learning rate of $53e-$ 5 and a linear warmup with $100$ steps. The batch size is set to $16$ for training, and we use a batch size of $1$ and a maximum of $128$ decoding steps during inference. Top-$p$ sampling is set to $0.9$, temperature $\tau=0.7$. The epoch is set to $5$. For preprocessing, we randomly split the dataset into train, validation and test set with a ratio of 8:1:1.
5.3. Baselines
@@
SOTA approaches mainly divides into extractive, abstractive and long-document-based methods, details are in Appendix A.3.
@@
Table 4. Results for the human evaluation of model-generated summaries by three intelligent analysts.
@@
(a) CTIS task (b) APS task
Figure 3. Input length analysis of two tasks.
Refer to caption Refer to caption
(a) CTIS task (b) APS task
Figure 4. Few shot experiments of the proposed two tasks.
5.4. Main Results
Automatic Evaluation.
We compare the performance of extractive models, abstractive models and long-document-based models on CTIS and APS tasks. As depicted in Table 3, extractive models perform the almost worst on both tasks, demonstrating that intelligence summarization in cybersecurity involves more than just extracting sentences. Interestingly, they obtain the best BERTScore, one possible reason is that these two models are based on BERT, leading to high BERTScore. Compared with MatchSum, abstractive model like BART-large (about 374M) achieves better result, about 10.12% and 6.18% gain on ROUGE-L on two tasks, which shows that the abstractive models are more suitable for these two tasks. Then, long-document-based model like Longformer (102M) obtains similar results compared with BART-base (about 121M). Notably, LLAMA2-7B (zero-shot) and ChatGPT do not attain further improvement on all the evaluation metrics, leaving noticeable room for future work on domain LLMs fine-tuning and few-shot learning on the benchmark. Finally, we find that performance is worse on the APS task, likely because the attack process exhibits properties like variability, concealment and complexity, making APS more challenging than general summarization. Enhancing models for APS task remains an important direction for future research.
Table 5. Case analysis on CTIS and APS tasks. Orange words mean the hallucination generated by the baseline. Blue words present the correct result. Red words indicate some of the key information that needs to be extracted.
@@ 5.4. Main Results
BART 1. The phishing email contains a malicious RTF file titled “face mask order.zip” 2. The RTF contains an exploit that triggers CVE-2017-11882 to execute arbitrary code. 3. The exploit extracts the OLE objects from RTF files. 4. The OLE object contains a VBScript script that executes an Equation Editor payload. 5. The Equated Editor payload connects to a command and control server to receive commands. 6. Command server sends commands to the C2 server to execute additional payloads. 7. Additional payloads are delivered to the infected …
@@
Human A/B Evaluation.
In addition to the automatic evaluation, human A/B evaluation (Peng et al., 2023) is conducted to validate the effectiveness of current SOTA models that BART-large and Longformer are compared. The results in Table 4 demonstrate the consistent conclusion with automatic evaluation. It can be seen that the summaries from BART-large are much more preferred than those of the baselines in terms of the three aspects. For example, compared with Longformer, BART is superior in terms of the coverage metric, indicating that the larger the number of parameters, the better the generated summaries. Besides, it is noted that BART does not significantly outperform Longformer in the relevance metric, and this is probably attributed to the powerful understanding ability of the PLMs, but BART still obtains decent improvements.
5.5. Input Length Analysis
@@
5.6. Few-shot Analysis
@@ A.1. Ethical Statement
A.2. Details of Regular Expression
@@
A.3. Details of Baseline Models
Several SOTA approaches are illustrated in this section for comparison. The models can be mainly divided into extraction-based, abstraction-based and long-document-based methods.
Extractive Model
BertSumExt (Liu and Lapata, 2019) is a neural extractive summarization model based on BERT (Devlin et al., 2019), which takes BERT as the sentence encoder and a Transformer layer as the document encoder. A classifier is utilized to perform sentence selection, then the model outputs the final summary. MatchSum (Zhong et al., 2020) formulates the extractive summarization task as a semantic text matching problem and develops a novel summary-level framework MatchSum, which generates a collection of possible candidate summaries from the output of BertSumEXT. The candidate that matches best with the document is selected as the final summarized version.
Abstractive Model
Abstractive summarization involves generating new phrases and sentences that convey the most important information from the source document, which uses semantic understanding and language generation technology to make a summarization. Hence, we fine-tune BART (Lewis et al., 2020) and T5 (Raffel et al., 2020) from the HuggingFace library.
BART (Lewis et al., 2020) is an encoder-decoder language model based on a sequence-to-sequence architecture, which achieves strong performance on a variety of NLP tasks like summarization, question answering and text generation after fine-tuning. T5 (Raffel et al., 2020) uses a standard Transformer encoder-decoder architecture, which converts all NLP tasks into a unified text-to-text format where the input and output are always text strings. T5 comes in several sizes including T5-Small, T5-Base, T5-Large, etc. In this paper, the setting is based on T5-Base.
Long-document-based Model
Long-document-based summarization refers to generating summaries for documents that are significantly longer than typical texts, e.g. 1,024 tokens, like Longformer(Beltagy et al., 2020), LLAMA2 (Touvron et al., 2023), etc.
Longformer(Beltagy et al., 2020) is an extension to the Transformer architecture, which leverages an attention mechanism whose computational complexity grows linearly as the length of the input sequence increases, making it easy to process thousands of tokens or more. LLAMA2 (Touvron et al., 2023) is a Transformer-based large language model, which is trained on massive text datasets to learn natural language patterns and generate human-like text. The context length of LLAMA2 can extend up to 4096 tokens, allowing it to understand and generate longer texts. In this paper, the 7 billion LLAMA2 is utilized to compare against fine-tuned models in zero-shot setting.

### v2

@@ 3.2 Data Statistics
4 Problem Formulation
5 Experiments
5.1 Evaluation Metrics
Following papers [15, 26], BERTScore [43] and ROUGE-$n$ (R-$n$) [20] are taken as evaluation metrics, which are widely used for evaluating the quality of summarization. BERTScore is a metric for evaluating text generation models, which measures the similarity between the generated text and reference text using contextual embeddings from pre-trained BERT. ROUGE-$n$ refers to the overlap of n-grams between the generated and reference summaries. Specifically, ROUGE-$1$ evaluates unigram overlap, ROUGE-$2$ measures bigram overlap, and ROUGE-$L$ calculates longest common subsequence overlap. These metrics compare matching units such as words and phrases between the generated and reference summaries, with higher scores indicating better summarization quality. Among the ROUGE metrics, ROUGE-$L$ generally corresponds best with human judgments of summary quality.
In previous studies, human evaluation is usually conducted by crowdsourcing workers who rate responses on a scale from 1 to 5 from the aspects of correctness, relevancy, etc. However, the criteria can vary widely between different individuals. Therefore, this study adopts the human A/B evaluation for a high inter-annotator agreement. Given the generated summaries of two models A and B, three analysts are prompted to go through an entire CTI report and choose the better one for each of the 80 randomly sub-sampled test instances. For objectivity, annotators include those with and without background knowledge (task-related). The final results are determined by majority voting. If the three annotators reach different conclusions, the fourth annotator will be brought in. We adopt the same human evaluation with paper [26]: 1) Factual Correctness: which summary can be supported by the source CTI report? 2) Relevance: which summary captures pertinent information relative to the CTI report? 3) Coverage: which summary contains the greatest coverage of relevant content in the CTI report?
Table 2: Performance of automatic evaluation on CTIS and APS tasks. LLAMA2 is in the zero-shot setting because of the lack of resources. The best results are highlighted in bold.
@@ 5.1 Evaluation Metrics
BertSumExt [22] 82.15 26.91 5.91 13.09 81.36 20.43 3.52 11.76
MatchSum [44] 83.91 39.31 13.39 20.60 83.75 33.76 9.34 18.91
@@
Transformer [41] 45.83 34.15 12.05 19.88 45.17 32.94 10.37 18.55
T5 [34] 70.48 43.75 16.58 28.06 69.21 42.41 14.35 27.32
BART-base [19] 70.36 44.45 16.26 27.24 69.24 42.48 14.30 26.23
BART-large [19] 71.21 47.11 18.32 29.20 70.41 45.76 16.88 29.03
@@
Longformer[3] 71.06 46.97 17.21 28.22 70.31 45.39 15.66 27.09
LLAMA2 [40] (zero-shot) 33.87 21.35 5.93 20.51 32.15 20.89 5.66 19.22
GPT-4o (zero-shot) 65.82 40.18 13.22 24.33 64.31 38.70 12.71 22.98
@@
BertSumExt [22] 81.38 20.67 3.06 11.87 81.38 20.35 3.46 11.73
MatchSum [44] 83.93 33.86 9.86 19.17 83.75 33.73 9.35 18.88
@@
Transformer [41] 40.43 28.22 7.86 16.50 39.15 27.03 6.21 15.48
T5 [34] 68.41 36.20 10.27 24.70 66.06 32.09 7.62 21.78
BART-base [19] 70.47 40.47 12.33 25.28 70.70 39.42 11.13 24.45
BART-large [19] 71.31 41.88 12.76 26.54 70.32 40.25 11.65 25.06
@@
Longformer[3] 70.98 41.35 11.77 25.36 70.41 39.91 9.78 23.93
LLAMA2 [40] (zero-shot) 21.06 15.68 2.23 13.14 20.38 14.66 2.02 12.28
GPT-4o (zero-shot) 48.77 32.06 9.11 18.69 45.39 30.55 8.61 17.75
5.2 Experimental Setting
@@
The implementation of baselines is based on the HuggingFace framework. The AdamW optimizer [24] with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ is used for training, with an initial learning rate of $53e-$ 5 and a linear warmup with $100$ steps. The batch size is set to $16$ for training, and we use a batch size of $1$ and a maximum of $128$ decoding steps during inference. Top-$p$ sampling is set to $0.9$, temperature $\tau=0.7$. The epoch is set to $5$. For preprocessing, we randomly split the dataset into train, validation and test set with a ratio of 8:1:1.
5.3 Baselines
@@
Several SOTA approaches are illustrated for comparison. The models can be mainly divided into extraction-based, abstraction-based and long-document-based methods.
Extractive Model
BertSumExt [22] is a neural extractive summarization model based on BERT [11], which takes BERT as the sentence encoder and a Transformer layer as the document encoder. A classifier is utilized to perform sentence selection, then the model outputs the final summary.
MatchSum [44] formulates the extractive summarization task as a semantic text matching problem and develops a novel summary-level framework MatchSum, which generates a collection of possible candidate summaries from the output of BertSumEXT. The candidate that matches best with the document is selected as the final summarized version.
Abstractive Model
Abstractive summarization involves generating new phrases and sentences that convey the most important information from the source document, which uses semantic understanding and language generation technology to make a summarization. Hence, we fine-tune BART [19] and T5 [34] from the HuggingFace library.
BART [19] is an encoder-decoder language model based on a sequence-to-sequence architecture, which achieves strong performance on a variety of NLP tasks like summarization, question answering and text generation after fine-tuning.
T5 [34] uses a standard Transformer encoder-decoder architecture, which converts all NLP tasks into a unified text-to-text format where the input and output are always text strings. T5 comes in several sizes including T5-Small, T5-Base, T5-Large, etc. In this paper, the setting is based on T5-Base.
Long-document-based Model
Long-document-based summarization refers to generating summaries for documents that are significantly longer than typical texts, e.g. 1,024 tokens, like Longformer[3], LLAMA2 [40], etc.
Longformer[3] is an extension to the Transformer architecture, which leverages an attention mechanism whose computational complexity grows linearly as the length of the input sequence increases, making it easy to process thousands of tokens or more. LLAMA2 [40] is a Transformer-based large language model, which is trained on massive text datasets to learn natural language patterns and generate human-like text. The context length of LLAMA2 can extend up to 4096 tokens, allowing it to understand and generate longer texts. In this paper, the 7 billion LLAMA2 is utilized to compare against fine-tuned models in zero-shot setting.
@@
Table 3: Results for the human evaluation of model-generated summaries by three intelligent analysts.
@@
5.4 Main Results
Automatic Evaluation.
We compare the performance of extractive models, abstractive models and long-document-based models on CTIS and APS tasks. As depicted in Table 2, extractive models perform the almost worst on both tasks, demonstrating that intelligence summarization in cybersecurity involves more than just extracting sentences. Interestingly, they obtain the best BERTScore, one possible reason is that these two models are based on BERT, leading to high BERTScore. Compared with MatchSum, abstractive model like BART-large (about 374M) achieves better result, about 10.12% and 6.18% gain on ROUGE-L on two tasks, which shows that the abstractive models are more suitable for these two tasks. Then, long-document-based model like Longformer (102M) obtains similar results compared with BART-base (about 121M). Notably, LLAMA2-7B and GPT-4o (zero-shot) do not attain further improvement on all the evaluation metrics, leaving noticeable room for future work on domain LLMs fine-tuning and few-shot learning on the benchmark. Finally, we find that performance is worse on the APS task, likely because the attack process exhibits properties like variability, concealment and complexity, making APS more challenging than general summarization. Enhancing models for APS task remains an important direction for future research.
@@ 5.4 Main Results
(a) CTIS task. (b) APS task.
Figure 3: Input length analysis of two tasks.
Table 4: Fleiss Kappa evaluation of CTISum dataset.
Score 1 2 3 4 5 Fleiss Kappa
Result 0 9 58 129 54 0.37
Score 1 and 2 3 4 and 5 Fleiss Kappa
Result 9 58 183 0.61
Human A/B Evaluation.
In addition to the automatic evaluation, human A/B evaluation [32] is conducted to validate the effectiveness of current SOTA models that BART-large and Longformer are compared. The results in Table 3 demonstrate the consistent conclusion with automatic evaluation. It can be seen that the summaries from BART-large are much more preferred than those of the baselines in terms of the three aspects. For example, compared with Longformer, BART is superior in terms of the coverage metric, indicating that the larger the number of parameters, the better the generated summaries. Besides, it is noted that BART does not significantly outperform Longformer in the relevance metric, and this is probably attributed to the powerful understanding ability of the PLMs, but BART still obtains decent improvements.
5.5 Fleiss Kappa Analysis
To verify the consistency of the final summary generation quality, we select 5 experts to execute a consistency evaluation of 50 sampled summaries using Fleiss Kappa, where the scores are rated from 1 to 5, with 1 very negative, 2 negative, 3 neutral, 4 positive, and 5 very positive. As shown in the Table 4, the final Fleiss Kappa score is 0.37, indicating relatively low agreement. This can be attributed to the fact that summary generation belongs to language generation task, where different experts tend to have subjective views on summary quality (e.g., examples that annotated as positive or very positive are inconsistent, but they are all good), which differs greatly from the objectivity of classification tasks. However, we also analyze the score distribution of all experts, finding that summaries rated 3 or above account for 96.4% (241/250) of the total, demonstrating that the quality of the annotated dataset is controllable and accurate. Moreover, we have conduct another round of Fleiss Kappa Analysis on the previous annotation scores from multiple raters. Specifically, we combine score 1 and 2 into one bin, keep score 3 as a separate bin, and combine score 4 and 5 into another bin. This result in a Fleiss Kappa score of 0.613184, which indicate the consistency evaluation.
Table 5: Case analysis on CTIS and APS tasks. Orange words mean the hallucination generated by the baseline. Blue words present the correct result. Red words indicate some of the key information that needs to be extracted.
@@ 5.5 Fleiss Kappa Analysis
BART 1. The phishing email contains a malicious RTF file titled “face mask order.zip" 2. The RTF contains an exploit that triggers CVE-2017-11882 to execute arbitrary code. 3. The exploit extracts the OLE objects from RTF files. 4. The OLE object contains a VBScript script that executes an Equation Editor payload. 5. The Equated Editor payload connects to a command and control server to receive commands. 6. Command server sends commands to the C2 server to execute additional payloads. 7. Additional payloads are delivered to the infected …
@@
5.6 Input Length Analysis
@@
Refer to caption Refer to caption
(a) CTIS task. (b) APS task.
Figure 4: Few shot experiments of the proposed two tasks.
5.7 Few-shot Analysis
@@ 6 Conclusion
Acknowledgments
We thank all anonymous reviewers for their constructive comments. This work is supported by the Zhongguancun Laboratory.
@@ A.1 Ethical Statement
A.2 Details of Regular Expression
@@
[…]